OpenStudy (anonymous):

Solve for a: 4 a over 9 = 100 over 45.

OpenStudy (anonymous):

\[\frac{ 4a }{ 9 } = \frac{ 100 }{ 45 }\]

OpenStudy (anonymous):

so u know about cross multiplication?

OpenStudy (anonymous):

No.

OpenStudy (anonymous):

ok the thing is, if \[\frac{ a }{ b } = \frac{ c }{ d } => a . d = c . b \]

OpenStudy (anonymous):

so can u try cross multiplication and make the equation ?

OpenStudy (anonymous):

108 and 900?

OpenStudy (anonymous):

45.4 = 180, not 108. :P and ur multiplying with 4a, so 45.4a = 180a, which is equal to 900. Hence \[180.a = 900.\]

OpenStudy (anonymous):

if u can find a now?

OpenStudy (anonymous):

Is it 5?

OpenStudy (anonymous):

yes it is 5 :)